1. Prepare a cost of production repor

1. Prepare a cost of production report, and identify the missing amounts for Work in Process—Roasting Department. If an amount is zero, enter in a zero "0". When computing cost per equivalent units, round to the nearest cent.

Fresh Mountain Coffee Company
Cost of Production Report-Roasting Department

For the Month Ended March 31, 2016
Answer #1

Working

Direct Material Cost per unit in Feb = 5700/1500 = $ 3.80

Direct Material Cost per unit in March = 3.90

Increase in Direct Material Cost per unit = 3.90 -3.80

Increase in Direct Material Cost per unit = 0.10

Conversion Cost per unit in Feb = (6150-5700)/(1500*30%) = $ 1

Conversion Cost per unit in March =0.80

Decrease in Conversion Cost per unit = 1 - 0.80

Decrease in Conversion Cost per unit = 0.20